Please help me solve this problem.
(x-5)^ 2/3 = 16
so it's (x-5) to the 2/3 power equals 16, but I don't know how to distribute the 2/3 to continue my equation

Answer by Alan3354(69443)

Please help me solve this problem.
(x-5)^ 2/3 = 16
--------
cube both sides
(x-5)^2 = 16^3 = 4096
----
Take the sq roots
x-5 = 64
x = 69
=======
x-5 = -64
x = -59